### v2

@@ 6.1.1 Proof of Theorem 4
We first show that TTC induces a proper core allocation. Suppose that TTC does not induce a proper core allocation. Then, there are $R\in\mathcal{R}^{N}$ and $S\subseteq N$ such that $S$ properly blocks $TTC(R)$ via an allocation $y\neq TTC(R)$. Let $x\equiv TTC(R)$. Let $\mathcal{C}\in\mathcal{C}_{1}(R)$ be a trading cycle formed at the first step under $TTC(R)$, and recall that $S_{\mathcal{C}}\subseteq N$ is the set of agents who are involved in $\mathcal{C}$. There are two cases. First, if an agent $i\in S_{\mathcal{C}}$ is also in $S$, then we have $y_{i}=x_{i}$ as $x_{i}$ is his most preferred object. Then, by our self-autarky condition (2), there is an agent $\widehat{i}\in S$ such that $x_{i}\in T_{\widehat{i}}(R)$. It implies that $\widehat{i}\in S_{\mathcal{C}}$; otherwise $i$ cannot receive $x_{i}$ via TTC. Consequently, we conclude that all agents in $S_{\mathcal{C}}$ are in $S$ and receive their TTC allocation at $y$ as well. This implies that no agent in $S$ can be strictly better off by receiving an object from $\{x_{i}\}_{i\in S_{\mathcal{C}}}$. Second, if no agent in $S_{\mathcal{C}}$ is in $S$, then by Claim 3, we see that for each agent in $S$, his trading opportunity set does not contain any object from $\{x_{i}\}_{i\in S_{\mathcal{C}}}$. In any case, we conclude that no agent in $S$ can be strictly better off by receiving an object from an object that is removed from the first step of TTC. Applying the argument to the second step and so on, we obtain that there is no agent $j\in S$ such that $y_{j}\mathbin{P_{j}}x_{j}$, leading to a contradiction.
@@ 6.2 Pairwise stability
For priority-based allocation problems, e.g., school choice problems, a standard stability notion is pairwise stability (Gale and Shapley, 1962): An allocation $x$ is pairwisely stable at $R$ if there is no agent-object pair $(i,o)$ such that $o\mathbin{P_{i}}x_{i}$ and $i\succ_{o}j$, where $x_{j}=o$. Clearly, this notion is not suitable here, as in our model, objects have no priorities over agents. Even if we adopt each center $c$’s priority for objects in $O_{c}$, it would still be problematic because $\succ_{c}$ is incomplete. One way to address this issue is to expand the priorities to be complete. A natural expansion is as follows: For each $c\in C$, each $o\in O_{c}$, each $i\in N_{c}$, and each $j\not\in N_{c}$, $i$ has higher expanded priority at $o$ than jj. [^18] Karakaya et al. (2019) propose an extension of TTC based on such expanded priorities, and Combe (2023) shows that this TTC extension is “minimal envy” among the set of individually rational, strategy-proof, and Pareto efficient mechanisms. However, this assumption will lead to an incompatibility with pair efficiency. To see it, let us revisit Example 1. Suppose that at $a$ and $b$, agent $1$ has higher priority than $3$, as $a,b\in O_{c},1\in N_{c}$, and $3\not\in N_{c}$. Similarly, $3$ has higher priority than $1$ and $2$ at $d$. Suppose that $2$ receives $d$: in this case, if $3$ receives $b$, then $3$ would block this allocation as $3$ has higher priority than $2$ at $d$; however, if $3$ receives $a$ then it means $1$ receives $b$. So, $1$ would block this allocation as $1$ has higher priority than $3$ at $a$. Thus, $2$ does not receive $d$. Since $2$ also has priority than $1$ at $a$, we conclude that it must be the case that $2$ receives $a$. Subsequently, $1$ receives $b$ and $3$ receives $d$, i.e., the allocation $y$. However, this allocation is not pair efficient: agents $2$ and $3$ can be better off by switching their allotments.
@@ 7.1 Length constraints
The primary application target of this paper is cross-program cooperation in kidney exchange. One might wonder if TTC is not applicable for kidney exchanges, as it allows trading cycles without length constraints (of course, in our model, each cycle contains at most $|C|$ agents). [^19] If $|C|$ is sufficiently small, we may not even meet length constraints. Note that in India, cyclic exchanges involving up to ten pairs have been successfully conducted (Kute et al., 2021). In practice, cycle length constraints are imposed due to the scarcity of medical resources: a two-way exchange requires the simultaneous availability of four operating rooms and associated personnel, while a three-way exchange requires six, and so on. However, in our model, the presence of multiple centers allows for integrating medical resources from different centers, making it feasible to conduct exchanges with longer cycles. [^20] There are also studies that examine cooperation across different kidney exchange programs without considering length constraints; see, e.g., Benedek et al. (2024).
@@ 7.2 Balancedness
Another restriction in our model is balancedness: For each center $c$, the inflow of agents coming from other centers to receive objects in $O_{c}$ is equal to the outflow of agents in $N_{c}$ who leave $c$ to receive objects in other centers. This restriction aligns with real-world scenarios in kidney exchanges: As in the classical kidney exchange model, assuming $|N_{c}|=|O_{C}|$ is based on the observation that, in most cases, each patient comes with a living kidney donor, often a family member. While this restriction seem to limit the applicability of this paper to other use cases, we argue this point by the following reason: Balancedness is an inviolable constraint in many practical allocation problems, such as shift reallocation and exchange programs (Dur and Ünver, 2019; Yu and Zhang, 2020; Manjunath and Westkamp, 2021; Kamada and Kojima, 2022). As pointed out by Dur and Ünver (2019), in such settings, every position and every worker must be matched, and hence the balancedness restriction is imposed. Thus, balancedness does not reduce the applicability of this paper. Of course, considering extensions involving imbalance or varying object capacities is an interesting open problem. However, such issues are beyond the scope of this paper, and we leave them for future research.
@@ A.1 Proof of Lemma 1
Let $f$ be procedurally fair. Let $R\in\mathcal{R}^{N}$ and $x\equiv f(R)$. Let $i\in N$, $c\equiv c(i)$. Suppose that $i$ has the $k$-highest priority at $c$, i.e., $r_{i}=k$. We see that at most k1k-1 objects in $O_{c}$ are not in $T_{i}(x)$. Consider the worst case that all the k1k-1 number of objects are more preferred than $o_{i}(R)$. In this case, $o_{i}(R)\in T_{i}(x)$. Therefore, $x_{i}\mathbin{R_{i}}o_{i}(R)$ holds, because at least agent $i$ can directly consume $o_{i}(R)$ from $T_{i}(x)$.
@@ A.2 Proof of Proposition 1
External fairness. By the definition of TTC, we know that each agent in a cycle depends on each of the others in that cycle (or agents in the previous cycles). Also we know that, if $i$ envies $j$, then $j$ receives his allotment at an earlier step than $i$. Let $\mathcal{C}$ be the cycle that includes $j$. So, $i\in N_{c(i)}$ and $j\not\in N_{c(i)}$. In $\mathcal{C}$, the object $i$ envies is pointing to an agent $k\in N_{c(i)}$ who is higher ranked than $i$ (i.e., $k\succ_{c(i)}i$) and who depends on $j$. Therefore, this envy is not externally justified.
@@ Appendix B Independence examples
Staregy-proofness: Example 5 is procedurally fair but is not strategy-proof.
Procedural fairness: A constant mechanism that always selects the same allocation is strategy-proof but is not procedurally fair.
Next, we show that procedural fairness is independent with other mentioned properties. By Theorem 2 and independence examples in Theorem 1, one can find that strategy-proofness / pair efficiency / internal fairness / external fairness cannot imply procedural fairness solely. Thus, we only show that procedural fairness does not imply any of the other properties.
@@
Consider $x\equiv(o_{3},o_{4},o_{2},o_{1})$. Note that $T_{1}(x)=\{o_{1},o_{2}\},T_{3}=\{o_{3}\},T_{4}(x)=\{o_{4}\}$, and $T_{2}(x)$ is either $\{o_{1}\}$ or $\{o_{2}\}$. One can see that agents $1$ and $2$ cannot be better off because they receives their most preferred objects. Also, agents $3$ and $4$ prefer any object in $\{o_{1},o_{2}\}$ over any object in $\{o_{3},o_{4}\}$. Thus, $x$ is fairly produced. However, $x$ is not pair efficient as agents $3$ and $4$ can be better off by switching their allotments.
@@
Consider $\equiv(o_{2},o_{3},o_{1}). xx$ is not internally fair as $1$ envies $2$ and $1\succ_{c_{1}}2$. However, $x$ is fairly produced. To see it, first note that $T_{1}(x)=\{o_{1},o_{2}\},T_{3}=\{o_{3}\}$, and $T_{2}(x)$ is either $\{o_{1}\}$ or $\{o_{2}\}$. So, while agent $1$ prefers $o_{2}$ to $o_{1}$, $1$ cannot be better off solely via consuming another object from $T_{1}(x)$. Also note that only agent $1$ does not receive his most preferred object. It means that agent $1$ cannot be better off by trading with other agents. Finally,
External fairness: Example 2 is procedurally fair but is not externally fair.

### v3

@@ 6.1.1 Proof of Theorem 4
We first show that TTC induces a proper core allocation. Suppose that TTC does not induce a proper core allocation. Then, there are $R\in\mathcal{R}^{N}$ and $S\subseteq N$ such that $S$ properly blocks $TTC(R)$ via an allocation $y\neq TTC(R)$. Let $x\equiv TTC(R)$. Let $\mathcal{C}\in\mathcal{C}_{1}(R)$ be a trading cycle formed at the first step under $TTC(R)$, and recall that $S_{\mathcal{C}}\subseteq N$ is the set of agents who are involved in $\mathcal{C}$. There are two cases. First, if an agent $i\in S_{\mathcal{C}}$ is also in $S$, then we have $y_{i}=x_{i}$ as $x_{i}$ is his most preferred object. Then, by our self-autarky condition (2), there is an agent $\widehat{i}\in S$ such that $x_{i}\in T_{\widehat{i}}(R)$. It implies that $\widehat{i}\in S_{\mathcal{C}}$; otherwise $i$ cannot receive $x_{i}$ via TTC. Generally, we find that for each object involved in $\mathcal{C}$ there is a unique agent whose trading opportunity set contains that object, namely, the agent in $S_{\mathcal{C}}$ to whom the object points. Together with our self-autarky condition (2), it implies that agent is also in $S$. Consequently, we conclude that $S_{\mathcal{C}}\subseteq S$, and all agents in $S_{\mathcal{C}}$ receive their TTC allocation at $y$ as well. This implies that no agent in $S$ can be strictly better off by receiving an object from $\{x_{i}\}_{i\in S_{\mathcal{C}}}$. Second, if no agent in $S_{\mathcal{C}}$ is in $S$, then by Claim 3, we see that for each agent in $S$, his trading opportunity set does not contain any object from $\{x_{i}\}_{i\in S_{\mathcal{C}}}$. In any case, we conclude that no agent in $S$ can be strictly better off by receiving an object from an object that is removed from the first step of TTC. Applying the argument to the second step and so on, we obtain that there is no agent $j\in S$ such that $y_{j}\mathbin{P_{j}}x_{j}$, leading to a contradiction.
@@ 6.2 Pairwise stability
For priority-based allocation problems, e.g., school choice problems, a standard stability notion is pairwise stability (Gale and Shapley, 1962): An allocation $x$ is pairwisely stable at $R$ if there is no agent-object pair $(i,o)$ such that $o\mathbin{P_{i}}x_{i}$ and $i\succ_{o}j$, where $x_{j}=o$. Clearly, this notion is not suitable here, as in our model, objects have no priorities over all agents. Even if we adopt each center $c$’s priority for objects in $O_{c}$, it would still be problematic because $\succ_{c}$ is incomplete. One way to address this issue is to expand the priorities to be complete. A natural expansion is as follows: For each $c\in C$, each $o\in O_{c}$, each $i\in N_{c}$, and each $j\not\in N_{c}$, $i$ has higher expanded priority at $o$ than jj. [^19] Karakaya et al. (2019) propose an extension of TTC based on such expanded priorities, and Combe (2023) shows that this TTC extension is “minimal envy” among the set of individually rational, strategy-proof, and Pareto efficient mechanisms. However, this assumption will lead to an incompatibility with pair efficiency. To see it, let us revisit Example 1. Suppose that at $a$ and $b$, agent $1$ has higher priority than $3$, as $a,b\in O_{c},1\in N_{c}$, and $3\not\in N_{c}$. Similarly, $3$ has higher priority than $1$ and $2$ at $d$. Suppose that $2$ receives $d$: in this case, if $3$ receives $b$, then $3$ would block this allocation as $3$ has higher priority than $2$ at $d$; however, if $3$ receives $a$ then it means $1$ receives $b$. So, $1$ would block this allocation as $1$ has higher priority than $3$ at $a$. Thus, $2$ does not receive $d$. Since $2$ also has priority than $1$ at $a$, we conclude that it must be the case that $2$ receives $a$. Subsequently, $1$ receives $b$ and $3$ receives $d$, i.e., the allocation $y$. However, this allocation is not pair efficient: agents $2$ and $3$ can be better off by switching their allotments.
@@ 7.1 Length constraints
The primary application target of this paper is cross-program cooperation in organ exchange programs, e.g., kidney and liver exchanges. One might wonder if TTC is not applicable for kidney exchanges, as it allows trading cycles without length constraints (of course, in our model, each cycle contains at most $|C|$ agents). [^20] If $|C|$ is sufficiently small, we may not even meet length constraints. Note that in India, cyclic exchanges involving up to ten pairs have been successfully conducted (Kute et al., 2021). In practice, cycle length constraints are imposed due to the scarcity of medical resources: a two-way exchange requires the simultaneous availability of four operating rooms and associated personnel, while a three-way exchange requires six, and so on. However, in our model, the presence of multiple centers allows for integrating medical resources from different centers, making it feasible to conduct exchanges with longer cycles. [^21] There are also studies that examine cooperation across different kidney exchange programs without considering length constraints; see, e.g., Benedek et al. (2024).
@@ 7.2 Balancedness
Another restriction in our model is balancedness: For each center $c$, the inflow of agents coming from other centers to receive objects in $O_{c}$ is equal to the outflow of agents in $N_{c}$ who leave $c$ to receive objects in other centers. This restriction aligns with real-world scenarios in kidney exchanges: As in the classical kidney exchange model, assuming $|N_{c}|=|O_{C}|$ is based on the observation that, in most cases, each patient comes with a living kidney donor, often a family member. [^22] For organ exchange environments with imbalance and capacity constraints, Ekici (2024b) proposes a solution by introducing fictitious agents and objects; the same construction can be applied here as well. While this restriction seem to limit the applicability of this paper to other use cases, we argue this point by the following reason: Balancedness is an inviolable constraint in many practical allocation problems, such as shift reallocation and exchange programs (Dur and Ünver, 2019; Yu and Zhang, 2020; Manjunath and Westkamp, 2021; Kamada and Kojima, 2025). As pointed out by Dur and Ünver (2019), in such settings, every position and every worker must be matched, and hence the balancedness restriction is imposed. Thus, balancedness does not reduce the applicability of this paper. Of course, considering extensions involving imbalance or varying object capacities is an interesting open problem. However, such issues are beyond the scope of this paper, and we leave them for future research.
@@ A.1 Proof of Lemma 1
Let $f$ be procedurally fair. Let $R\in\mathcal{R}^{N}$ and $x\equiv f(R)$. Let $i\in N$, $c\equiv c(i)$. Suppose that $i$ has the $k$-highest priority at $c$, i.e., $r_{i}=k$. We see that at most k1k-1 objects in $O_{c}$ are not in $T_{i}(R)$. Consider the worst case that all the k1k-1 number of objects are more preferred than $o_{i}(R)$. In this case, $o_{i}(R)\in T_{i}(R)$. Therefore, $x_{i}\mathbin{R_{i}}o_{i}(R)$ holds, because at least agent $i$ can directly consume $o_{i}(R)$ from $T_{i}(R)$.
@@ A.2 Proof of Proposition 1
External fairness. By the definition of TTC, we know that each agent in a cycle depends on each of the others in that cycle (or agents in the previous cycles). Also we know that, if $i$ envies $j$, then $j$ receives his allotment, denoted by $o$, at an earlier step than $i$. Let $\mathcal{C}$ be the cycle that includes $j$. So, $i\in N_{c(i)}$ and $j\not\in N_{c(i)}$, and recall that $o\in O_{c(i)}$. In $\mathcal{C}$, $o$ is pointing to an agent $k\in N_{c(i)}$ who is higher ranked than $i$ (i.e., $k\succ_{c(i)}i$) and who depends on $j$. Therefore, this envy is not externally justified.
@@ Appendix B Independence examples
Staregy-proofness: Consider a modification of Example 5 as follows.
Let $\widetilde{R}$ be such that
$\widetilde{R}_{1}:o_{2},o_{3},o_{1},o_{4};$
$\widetilde{R}_{2}:o_{1},o_{3},o_{2},o_{4};$
$\widetilde{R}_{3}:o_{1},o_{2},o_{3},o_{4}.$
$\widetilde{R}_{4}:\ldots$
Let $f$ be such that $f(\widetilde{R})=(o_{2},o_{3},o_{1},o_{4})$ and, for each $R\neq\widetilde{R}$, $f(R)=TTC(R)$. $f$ is procedurally fair but is not strategy-proof. $\diamond$
Procedural fairness: A constant mechanism that always selects the same allocation is strategy-proof but is not procedurally fair. $\diamond$
Next, we show that procedural fairness is independent with other mentioned properties. By Theorem 2 and independence examples in Theorem 1, one can find that pair efficiency / internal fairness / external fairness cannot imply procedural fairness solely. Thus, we only show that procedural fairness does not imply any of the other properties.
@@
Consider $x\equiv(o_{3},o_{4},o_{2},o_{1})$. Note that $T_{1}(R)=\{o_{1},o_{2}\},T_{3}(R)=\{o_{3}\},T_{4}(R)=\{o_{4}\}$, and $T_{2}(R)=\{o_{2}\}$. One can see that agents $1$ and $2$ cannot be better off because they receives their most preferred objects. Also, agents $3$ and $4$ prefer any object in $\{o_{1},o_{2}\}$ over any object in $\{o_{3},o_{4}\}$. Thus, $x$ is fairly produced. However, $x$ is not pair efficient as agents $3$ and $4$ can be better off by switching their allotments. $\diamond$
@@
Consider $\equiv(o_{2},o_{3},o_{1}). xx$ is not internally fair as $1$ envies $2$ and $1\succ_{c_{1}}2$. However, $x$ is fairly produced. To see it, first note that $T_{1}(x)=\{o_{1},o_{2}\},T_{3}=\{o_{3}\}$, and $T_{2}(x)$ is either $\{o_{1}\}$ or $\{o_{2}\}$. So, while agent $1$ prefers $o_{2}$ to $o_{1}$, $1$ cannot be better off solely via consuming another object from $T_{1}(x)$. Also note that only agent $1$ does not receive his most preferred object. It means that agent $1$ cannot be better off by trading with other agents. $\diamond$
External fairness: Consider a modification of Serial-QR mechanism in Example 2, where $Y(0)$ is adjusted to include all fairly produced allocations. This mechanism is procedurally fair but is not externally fair. $\diamond$